# Boundary Effects and Oxygen Deficiency-Driven Pattern Transitions in Algal Bioconvection

### v1

@@ Abstract
Suspensions of motile microorganisms can spontaneously give rise to large-scale fluid motion, known as bioconvection, which is characterized by dense, cell-rich downwelling plumes interspersed with broad upwelling regions. In this study, we investigate bioconvection in shallow suspensions of Chlamydomonas reinhardtii cells confined within spiral-shaped boundaries, combining detailed experimental observations with three-dimensional simulations. Under open liquid–air interface conditions, cells accumulate near the surface due to negative gravitaxis, forming spiral-shaped density patterns that subsequently fragment into lattice-like structures and give rise to downwelling plumes. Space–time analyses reveal coherent rotational dynamics, with inward-moving patterns near the spiral core and bidirectional motion farther from the center. Introducing confinement by sealing the top boundary with an air-impermeable transparent wall triggers striking transitions in the bioconvection patterns, driven by oxygen depletion: initially stable structures reorganize into new patterns with reduced characteristic wavelengths. Further eliminating the fluid gap above the spiral boundary extends the duration of this transitional phase, highlighting the essential role of fluid-mediated cell-to-cell communication in coordinating collective behavior. Complementary 3D simulations, based on the incompressible Navier–Stokes equations and incorporating negative buoyancy and active stress from swimming cells, capture the initial pattern formation and its subsequent instability, reproducing the fragmentation of spiral-shaped accumulations into downwelling plumes and the emergence of strong vortical flows—nearly an order of magnitude faster than individual cell swimming speeds. However, these models do not capture the oxygen-driven pattern transitions observed experimentally. Our findings reveal that confinement geometry, oxygen dynamics, and metabolic transitions critically govern bioconvection pattern evolution, offering new strategies to control microbial self-organization and flow through environmental and geometric design.
  • These authors contributed equally to this work.
Keywords: Chlamydomonas reinhardtii, Bioconvection, Micro-swimmers, Hydrodynamic instabilities, Aerobic to anaerobic metabolism transition.
@@ 1 Introduction
Microswimmers such as Escherichia coli, Bacillus subtilis, and the freshwater alga Chlamydomonas reinhardtii generate characteristic fluid flows through the coordinated beating of their flagella and body movements [1, 2, 3, 4, 5, 6, 7]. Depending on their flagellar arrangement, these microorganisms are typically categorized as either ”pushers” or ”pullers” [8, 9, 4, 10, 11]. In pusher-type swimmers, such as E. coli, the thrust generated by flagellar motion propels fluid outward along the swimming axis and draws it inward from the sides, creating extensile flow fields. Conversely, puller-type swimmers, such as Chlamydomonas reinhardtii, use anterior flagella to pull themselves forward, drawing fluid inward along the swimming axis and pushing it outward laterally, producing contractile flow fields. These microscale hydrodynamic disturbances significantly influence the motion of nearby passive particles and neighboring microorganisms, leading to complex collective behaviors, especially in dense suspensions. In particular, when large populations of motile microorganisms interact through the flows they generate, they can spontaneously organize into coherent large-scale structures, producing macroscopic flows through a process known as bioconvection. Unlike thermally driven convection, bioconvection arises solely from the active motility and orientation biases of the cells [9, 12, 13, 14, 15, 16, 17], and is characterized by the emergence of dense, cell-rich downwelling plumes interspersed with broad regions of upwelling fluid.
Bioconvection is commonly observed in shallow laboratory suspensions, where cells swim upward on average due to directional responses known as ”taxes,” including gravitaxis (orientation against gravity), gyrotaxis (a balance between gravitational and viscous torques), phototaxis (response to light), and chemotaxis (response to chemical gradients) [18, 19, 15, 19, 9, 12, 20, 21, 22]. Gravitaxis, in particular, is typically a passive orientation mechanism resulting from physical asymmetry within the cells, such as bottom-heavy mass distribution [12, 23, 24]. Cells swimming upward due to negative gravitaxis accumulate near the upper suspension boundary, creating an unstable density gradient that triggers convective instabilities and leads to the emergence of bioconvection patterns [13, 9, 15]. The fundamental mechanisms underlying bioconvection have been extensively modeled through continuum approaches [16, 25, 8], incorporating momentum and mass conservation equations coupled with equations governing the orientation and distribution of swimming cells. Early work by Kessler [12, 24] and later developments by Pedley and Kessler [9] established the theoretical framework for understanding how gyrotaxis and bottom-heaviness drive these large-scale convective phenomena.
@@
Recent studies have demonstrated that cell geometry, fluid depth, and boundary conditions critically influence the formation and stability of bioconvection patterns [18, 26, 27, 28, 29, 30]. While extensive research has explored bioconvection in simple geometries, the effects of more intricate internal structures, such as spiral-shaped boundaries, remain largely unexplored. Spiral geometries are particularly intriguing because they introduce geometrical singularities—regions of sharp curvature and varying local confinement—that can profoundly modulate cell accumulation, flow patterns, and instabilities. Additionally, the presence or absence of an open fluid–air interface significantly impacts bioconvection dynamics, primarily by regulating oxygen exchange between the suspension and the environment. Recent work by Fragkopoulos et al.[31] has shown that metabolic activity—closely linked to both light intensity and oxygen availability—directly modulates microbial motility and fundamentally governs the emergence and stability of coherent flow structures. These findings underscore the pivotal role of metabolic constraints, associated with light and oxygen dynamics, in shaping the spatiotemporal evolution of bioconvection patterns[13, 19, 31, 32].
@@
In this study, we investigate the formation and evolution of bioconvection patterns in suspensions of Chlamydomonas reinhardtii confined within spiral-shaped boundaries, under controlled red-light illumination to minimize phototactic effects. By systematically varying the boundary conditions—from open liquid–air interfaces to fully confined, air-impermeable geometries—we investigate how geometric confinement and oxygen depletion shape the spatiotemporal organization and characteristic wavelengths of bioconvection patterns. Through a combination of detailed experimental approaches—including top-view imaging, space–time analyses, and Fourier transforms—we show that confinement and oxygen depletion induce pronounced transitions in pattern morphology, resulting in the spontaneous reorganization of bioconvection patterns under oxygen-limited conditions. Complementary three-dimensional numerical simulations, based on the incompressible Navier–Stokes equations coupled with active-swimmer models, successfully reproduce the primary stages of pattern formation and hydrodynamic instability. However, these simulations do not capture the oxygen-dependent transitions observed experimentally. Together, our findings demonstrate that variations in confinement geometry and metabolic conditions can significantly alter collective microbial behavior, underscoring a strong interplay between microscale motility, environmental constraints, and emergent fluid flows.
@@ 2.1 Cell culture
[…]
@@ 2.2 Experimental setup
To investigate the influence of curved boundaries on bioconvection patterns in CR suspensions, we fabricated molds from acrylic (PMMA)—a material impermeable to air—with different internal geometries, including spiral-shaped and circular designs. Each design was examined under either open (aerobic) or sealed (anaerobic) top boundary conditions to systematically probe the effects of oxygen availability. The experimental setup was carefully leveled and filled with a thin layer of cell suspension (typically 5 mm deep) at a concentration of approximately $10^{7}$ cells/mL. All experiments were conducted inside a dark enclosure to eliminate phototactic responses triggered by ambient light. To further suppress phototactic effects during imaging and pattern development, the suspension was illuminated exclusively from above using a red LED light source centered at 625 nm (LUMIMAX, LSR24-R), with an optical diffuser positioned between the light source and the sample to ensure uniform illumination. Previous studies have demonstrated that the phototactic sensitivity of CR cells declines sharply at longer red wavelengths [35, 36, 37]. Therefore, illumination at red light effectively eliminates phototactic steering, ensuring that the observed bioconvection patterns primarily result from hydrodynamic instabilities and gravitactic behavior.
@@ 2.3 Image capture and analysis
For each experiment, over 1000 images were acquired at 1-second intervals using a dark-field imaging setup. In experiments requiring extended observation, patterns were recorded continuously for up to two hours at 1fps. The red LED illumination intensity, measured at the location of the PMMA mold, was maintained at approximately $5\times 10^{17}$ photons m-2 s-1. In the raw dark-field images, regions of high cell density appeared dark, whereas regions of low cell density appeared bright. To quantify the dominant wavelength of the emerging bioconvection patterns, Fourier analysis was performed on space–time plots using custom MATLAB scripts. Prior to analysis, image preprocessing was performed using FIJI. Initially, the images were inverted in intensity, and the mean image was subtracted from the stack to enhance the signal-to-noise ratio. In the resulting 16-bit post-processed images, regions of high cell density appeared bright (visualized in pseudo-green), whereas regions of low density appeared dark.
@@ 2.4 Model
Bioconvection, the spontaneous formation of large-scale patterns in suspensions of swimming microorganisms, arises from the interplay between individual cell motility, fluid mechanics, and external fields such as gravity [25, 38, 15, 14, 39, 40, 8]. Most theoretical models of bioconvection aim to link macroscopic flow fields to microscopic swimming behaviors, often under the assumption that direct cell-cell interactions are negligible due to the dilute nature of the suspension. In these systems, the Reynolds number associated with individual swimmers is extremely low (around $10^{-3}$), reflecting viscous-dominated motion at the microscale. However, the collective flows generated by bioconvection can reach much higher Reynolds numbers, often on the order of $10$, highlighting a dramatic amplification of fluid motion through coordinated swimming [28, 8, 14, 4].
Table 1: Parameters used in simulations of bioconvection patterns.
Mean concentration $\bar{c}$                            $4\times 10^{6}$ cells/ml
Fluid depth $L$                            5 mm
Fluid density $\rho_{0}$                            997 kg/m3
Cell density $\rho$                            1050 kg/m3
Specific gravity $\Delta\rho/\rho=\frac{\rho-\rho_{0}}{\rho}$           0.05
Cell radius $a$                            $5\times 10^{-6}$ m
Center of gravity offset $h$                            $10^{-5}$ cm
Volume per cell $\beta$                            $5\times 10^{-16}$ m3
Swimming speed $W_{c}$                         $10^{-4}$ m s-1
Sedimentation speed $W_{s}$                         $6\times 10^{-6}$ m s-1
Diffusivity of cells $D$                           $5\times 10^{-8}$ m2 s-1
Kinematic viscosity $\nu$                            $8.9\times 10^{-7}$ m2 s-1
Gyrotactic reorientation parameter $B$                            1 s
Scaled swimming speed $V_{c}=W_{c}L/D$           10
Schmidt number $S_{c}=\nu/D$               17.85
Gyrotaxis number $G=BD/L^{2}$           $2\times 10^{-3}$
Pseudo-Rayleigh number $R_{a}=\frac{\bar{c}\beta\Delta\rho gL^{3}}{(\rho\nu D)}$           2920
We model the collective dynamics of a dilute suspension of CR cells using a continuum framework based on earlier studies [25, 8, 16]. The model couples the incompressible Navier–Stokes equations with a conservation equation for the cell concentration and an orientation dynamics equation accounting for gyrotaxis:
@@
where $\bm{U}$ is the fluid velocity, $\bm{\omega}=\nabla\times\bm{U}$ is the vorticity field, $\mu$ is the fluid viscosity, $\rho_{0}$ is the fluid density, $p$ is the pressure, $c$ is the local cell concentration, and $\bm{p}$ denotes the local swimming direction of the cells. The stress tensor $\bm{E}=(\nabla\bm{U}+\nabla\bm{U}^{T})/2$ represents the rate of strain, and $\bm{\Sigma}$ is the active stress generated by the swimming cells, modeled as $\bm{\Sigma}=cS(\langle\bm{p}\bm{p}\rangle-\bm{I}/3)$ [41, 8, 16, 4]. Here $\bm{I}$ is the unity tensor, $S$ is the stresslet strength and is positive for CR cells which are ’pullers’ and negative for sperm cells which are pushers. The buoyancy term accounts for the density difference $\Delta\rho=\rho-\rho_{0}$ between cells and the surrounding fluid, and the Boussinesq approximation is employed [25, 16, 14]. Given that the characteristic timescale of bioconvection is considerably shorter than that of cellular growth, the conservation of the cell population is described by Eq. 3. This equation incorporates advection by both the fluid velocity and the active swimming motion at speed $W_{c}$, sedimentation under gravity at velocity $W_{s}$, and random motility modeled as an effective diffusion process with diffusivity $D$ [16, 8, 28]. Orientation dynamics (Eq. 4) describes the balance between gravitational torque, which favors upward orientation, and viscous torque arising from fluid shear, resulting in gyrotactic reorientation over a characteristic timescale given by $B=\alpha_{\perp}/2hmg,$ where $\alpha_{\perp}=8\pi\mu a^{3}$ is the rotational drag coefficient for a spherical cell of radius $a$, $h$ is the center-of-gravity offset within the cell, $m$ is the cell mass, and $g$ is the gravitational acceleration [8, 12]. This framework assumes the suspension is sufficiently dilute such that cell-cell hydrodynamic or steric interactions are negligible. Each cell is modeled as a sphere of volume $\beta$, swimming along its axis $\bm{p}$. Table 1 summarizes the parameters used in the numerical simulations, selected based on experimental measurements of CR cells and previous bioconvection studies [15, 4]. Numerical solutions of the coupled system were obtained using COMSOL Multiphysics (version 6.3), enabling us to reproduce key features of the experimentally observed bioconvection patterns.

### v3

@@ Abstract
Suspensions of motile microorganisms can spontaneously form large-scale fluid motion, known as bioconvection, characterized by dense downwelling plumes separated by broad upwelling regions. In this study, we investigate bioconvection in shallow suspensions of Chlamydomonas reinhardtii confined within spiral-shaped boundaries, combining detailed experiments with three-dimensional simulations. Under open liquid–air interfaces, cells accumulate near the surface via negative gravitaxis, generating spiral-shaped density patterns that subsequently fragment into lattice-like clusters, leading to plume formation. Space–time analyses demonstrate coherent rotational dynamics, with predominantly inward-directed motion near the spiral core and bidirectional motion further out. Introducing confinement by sealing the upper boundary with an air-impermeable wall triggers dramatic pattern transitions due to oxygen depletion: initially stable arrangements reorganize into new structures with significantly reduced wavelengths. Complementary numerical simulations, based on incompressible Navier–Stokes equations incorporating negative buoyancy and active swimmer stress, successfully replicate initial pattern formation, subsequent instability, fragmentation into plumes, and emergence of strong vortical flows—nearly an order of magnitude faster than individual cell swimming. However, these models do not capture oxygen depletion-driven transitions observed experimentally. Our results highlight that geometric confinement, oxygen availability, and metabolic transitions critically regulate bioconvection dynamics, offering novel strategies for controlling microbial self-organization and fluid transport.
footnotetext: ${}^{\text{a}}$ew York University Abu Dhabi, Abu Dhabi, United Arab Emirates. ${}^{\text{b}}$ Department of Physics, Lewis & Clark College, Portland, Oregon, USA.
E-mail: azam.gholami@nyu.edu
44footnotetext: These authors contributed equally to this work.footnotetext: † Electronic Supplementary Information (ESI) available: See DOI: 00.0000/$00$.
@@ 1 Introduction
Microswimmers such as Escherichia coli, Bacillus subtilis, and the freshwater alga Chlamydomonas reinhardtii generate characteristic fluid flows through the coordinated beating of their flagella and body movements 1, 2, 3, 4, 5, 6, 7. Depending on their flagellar arrangement, these microorganisms are typically categorized as either "pushers" or "pullers" 8, 9, 4, 10, 11. In pusher-type swimmers, such as E. coli, the thrust generated by flagellar motion propels fluid outward along the swimming axis and draws it inward from the sides, creating extensile flow fields. Conversely, puller-type swimmers, such as Chlamydomonas reinhardtii, use anterior flagella to pull themselves forward, drawing fluid inward along the swimming axis and pushing it outward laterally, producing contractile flow fields. These microscale hydrodynamic disturbances significantly influence the motion of nearby passive particles and neighboring microorganisms, leading to complex collective behaviors, especially in dense suspensions. In particular, when large populations of motile microorganisms interact through the flows they generate, they can spontaneously organize into coherent large-scale structures, producing macroscopic flows through a process known as bioconvection. Unlike thermally driven convection, bioconvection arises solely from the active motility and orientation biases of the cells 9, 12, 13, 14, 15, 16, 17, and is characterized by the emergence of dense, cell-rich downwelling plumes interspersed with broad regions of upwelling fluid.
[…]
@@
Recent studies have demonstrated that cell geometry, fluid depth, and boundary conditions critically influence the formation and stability of bioconvection patterns 18, 26, 27, 28, 29, 30. While extensive research has explored bioconvection in simple geometries, the effects of more intricate internal structures, such as spiral-shaped boundaries, remain largely unexplored. Spiral geometries are particularly intriguing because they introduce geometrical singularities—regions of sharp curvature and varying local confinement—that can profoundly modulate cell accumulation, flow patterns, and instabilities. Additionally, the presence or absence of an open fluid–air interface significantly impacts bioconvection dynamics, primarily by regulating oxygen exchange between the suspension and the environment. Recent work by Fragkopoulos et al.31 has shown that metabolic activity—closely linked to both light intensity and oxygen availability—directly modulates microbial motility and fundamentally governs the emergence and stability of coherent flow structures. These findings underscore the pivotal role of metabolic constraints, associated with light and oxygen dynamics, in shaping the spatiotemporal evolution of bioconvection patterns13, 19, 31, 32.
@@
In this study, we investigate the formation and evolution of bioconvection patterns in suspensions of Chlamydomonas reinhardtii confined within spiral-shaped boundaries, under controlled red-light illumination to minimize phototactic effects. By systematically varying the boundary conditions—from open liquid–air interfaces to fully confined, air-impermeable geometries—we investigate how geometric confinement and oxygen depletion shape the spatio-temporal organization and characteristic wavelengths of bioconvection patterns. Through a combination of detailed experimental approaches—including top-view imaging, space–time analyses, and Fourier transforms—we show that confinement and oxygen depletion induce pronounced transitions in pattern morphology, resulting in the spontaneous reorganization of bioconvection patterns under oxygen-limited conditions. Complementary three-dimensional numerical simulations, based on the incompressible Navier–Stokes equations coupled with active-swimmer models, successfully reproduce the primary stages of pattern formation and hydrodynamic instability. However, these simulations do not capture the oxygen-dependent transitions observed experimentally. Together, our findings demonstrate that variations in confinement geometry and metabolic conditions can significantly alter collective microbial behavior, underscoring a strong interplay between microscale motility, environmental constraints, and emergent fluid flows.
@@ 2.1 Cell culture
Chlamydomonas reinhardtii (CR) cultures were prepared following established protocols optimized to maintain high cell motility and physiological activity 33, 34. Wild-type CR cells (strain CC-409), obtained from the Algae Culture Collection at the University of Minnesota (USA), were initially cultured on agar plates containing Tris-Acetate-Phosphate (TAP) medium. The cultures were maintained under controlled illumination in a Memmert IPP260ecoplus incubator. To prepare liquid cultures, a small volume of cells from agar plates was inoculated into 500 mL Erlenmeyer flasks containing TAP medium and incubated at 24 C on an orbital shaker (INFORS HT Multitron) at 80 rpm under continuous illumination with warm white LEDs. Vegetative cells were harvested during the logarithmic growth phase, typically on the third or fourth day after inoculation. Cell densities were measured using an automated cell counter (PerkinElmer Cellometer) and used at the final concentration of approximately 4$\times$106 cells/mL.
@@ 2.2 Experimental setup
To investigate the influence of curved boundaries on bioconvection patterns in CR suspensions, we fabricated molds from acrylic (PMMA)—a material impermeable to air—with different internal geometries, including spiral-shaped and circular designs. Each design was examined under either open (aerobic) or sealed (anaerobic) top boundary conditions to systematically probe the effects of oxygen availability. The experimental setup was carefully leveled and filled with a thin layer of cell suspension (typically 5 mm deep) at a concentration of approximately 4$\times$106 cells/mL. All experiments were conducted inside a dark enclosure to eliminate phototactic responses triggered by ambient light. To further suppress phototactic effects during imaging and pattern development, the suspension was illuminated exclusively from above using a red LED light source centered at 625 nm (LUMIMAX, LSR24-R), with an optical diffuser positioned between the light source and the sample to ensure uniform illumination. Previous studies have demonstrated that the phototactic sensitivity of CR cells declines sharply at longer red wavelengths 35, 36, 37. Therefore, illumination at red light effectively eliminates phototactic steering, ensuring that the observed bioconvection patterns primarily result from hydrodynamic instabilities and gravitactic behavior.
@@ 2.3 Image capture and analysis
For each experiment, more than 1,000 top-view images were captured at 1-second intervals with an 8-bit depth. In experiments requiring longer observation, images were recorded continuously for up to two hours at a frame rate of 1 fps. The red LED illumination intensity, measured at the location of the PMMA mold, was maintained at approximately $5\times 10^{17}$ photons m-2 s-1. In the raw images, regions of high cell density appeared dark, whereas regions of low cell density appeared bright. To quantify the dominant wavelength of the emerging bioconvection patterns, Fourier analysis was performed on space–time plots using custom MATLAB scripts. Prior to analysis, image preprocessing was performed using FIJI. Initially, the images were inverted in intensity, and the mean image was subtracted from the stack to enhance the signal-to-noise ratio. In the resulting post-processed images, regions of high cell density appeared bright (visualized in pseudo-green), whereas regions of low density appeared dark.
@@ 2.4 Model
Bioconvection, the spontaneous formation of large-scale patterns in suspensions of swimming microorganisms, arises from the interplay between individual cell motility, fluid mechanics, and external fields such as gravity 25, 38, 15, 14, 39, 40, 8. Most theoretical models of bioconvection aim to link macroscopic flow fields to microscopic swimming behaviors, often under the assumption that direct cell-cell interactions are negligible due to the dilute nature of the suspension. In these systems, the Reynolds number associated with individual swimmers is extremely low (around $10^{-3}$), reflecting viscous-dominated motion at the microscale. However, the collective flows generated by bioconvection can reach much higher Reynolds numbers, often on the order of $10$, highlighting a dramatic amplification of fluid motion through coordinated swimming 28, 8, 14, 4.
We model the collective dynamics of a dilute suspension of CR cells using a continuum framework based on earlier studies 25, 8, 16. The model couples the incompressible Navier–Stokes equations with a conservation equation for the cell concentration and an orientation dynamics equation accounting for gyrotaxis:
@@
where $\bm{U}$ is the fluid velocity, $\bm{\omega}=\nabla\times\bm{U}$ is the vorticity field, $\mu$ is the fluid viscosity, $\rho_{0}$ is the fluid density, $P$ is the pressure, $c$ is the local cell concentration, and $\bm{p}$ denotes the local swimming direction of the cells. The stress tensor $\bm{E}=(\nabla\bm{U}+\nabla\bm{U}^{T})/2$ represents the rate of strain, and $\bm{\Sigma}$ is the active stress generated by the swimming cells, modeled as $\bm{\Sigma}=cS(\langle\bm{p}\bm{p}\rangle-\bm{I}/3)$ 41, 8, 16, 4, 42. Here $\bm{I}$ is the unity tensor, $S$ is the stresslet strength and is positive for CR cells which are ’pullers’ and negative for sperm cells which are pushers. The buoyancy term accounts for the density difference $\Delta\rho=\rho-\rho_{0}$ between cells and the surrounding fluid, and the Boussinesq approximation is employed 25, 16, 14. Given that the characteristic timescale of bioconvection is considerably shorter than that of cellular growth, the conservation of the cell population is described by Eq. 3. This equation incorporates advection by both the fluid velocity and the active swimming motion at speed $W_{c}$, sedimentation under gravity at velocity $W_{s}$, and random motility modeled as an effective diffusion process with diffusivity $D$ 16, 8, 28. Orientation dynamics (Eq. 4) describes the balance between gravitational torque, which favors upward orientation, and viscous torque arising from fluid shear, resulting in gyrotactic reorientation over a characteristic timescale given by $B=\alpha_{\perp}/2hmg,$ where $\alpha_{\perp}=8\pi\mu a^{3}$ is the rotational drag coefficient for a spherical cell of radius $a$, $h$ is the center-of-gravity offset within the cell, $m$ is the cell mass, and $g$ is the gravitational acceleration 8, 12. This framework assumes the suspension is sufficiently dilute such that cell-cell hydrodynamic or steric interactions are negligible. Each cell is modeled as a sphere of volume $\beta$, swimming along its axis $\bm{p}$. Table 1 summarizes the parameters used in the numerical simulations, selected based on experimental measurements of CR cells and previous bioconvection studies 15, 4. Numerical solutions of the coupled system were obtained using COMSOL Multiphysics (version 6.3), enabling us to reproduce some of the key features of the experimentally observed bioconvection patterns.
Table 1: Parameters used in simulations of bioconvection patterns.
Mean concentration, $\bar{c}$ $4\times 10^{6}\,\mathrm{cells/mL}$
Fluid depth, $L$ $5\,\mathrm{mm}$
Fluid density, $\rho_{0}$ $997\,\mathrm{kg\,m^{-3}}$
Cell density, $\rho$ $1050\,\mathrm{kg\,m^{-3}}$
Cell radius, $a$ $5\times 10^{-6}\,\mathrm{m}$
Center of gravity offset, $h$ $10^{-5}\,\mathrm{cm}$
Volume per cell, $\beta$ $5\times 10^{-16}\,\mathrm{m^{3}}$
Swimming speed, $W_{c}$ $10^{-4}\,\mathrm{m\,s^{-1}}$
Sedimentation speed, $W_{s}$ $6\times 10^{-6}\,\mathrm{m\,s^{-1}}$
Cell diffusivity, $D$ $5\times 10^{-8}\,\mathrm{m^{2}\,s^{-1}}$
Kinematic viscosity, $\nu$ $8.9\times 10^{-7}\,\mathrm{m^{2}\,s^{-1}}$
Gyrotactic parameter, $B$ $1\,\mathrm{s}$
Scaled swimming speed, $V_{c}=W_{c}L/D$ $10$
Schmidt number, $S_{c}=\nu/D$ $17.85$
Gyrotaxis number, $G=BD/L^{2}$ $2\times 10^{-3}$
Pseudo-Rayleigh number, $R_{a}=\dfrac{\bar{c}\,\beta\,\Delta\rho\,g\,L^{3}}{\rho\,\nu\,D}$ $2920$